# Extended thermodynamic analysis of a charged Horndeski black hole

### v1

@@ 1 Introduction
It is known that thermodynamic description of black holes is universal and can be applicable to black holes found in beyond Einstein gravity. However, it was found that the Wald’s entropy formula and /or the first law of thermodynamics does not work for the CHB [8]. To achieve the first law with the ordinary Hawking temperature, they introduced a new term $\Phi_{\phi}$ of the chemical potential and a conserved scalar charge $Q_{\phi}$. This derivation of the first law was based on the Wald’s entropy ($S_{W}=\pi(r_{+}^{2}-q^{2}/2)$) from the Noether’s theorem [9, 10] and the Hawking temperature ($T_{\kappa}=\kappa/2\pi$) obtained from the surface gravity. However, Noether’s theorem and Wald formula have ambiguities for the Einstein-Horndeski-Maxwell theory. These were fixed by deriving the new temperature and introducing the Bekenstein-Hawking entropy [11], leading to the standard first law of thermodynamics. Also, this case satisfies the Smarr formula of its integral form. We stress that this happened to the Bardeen regular black hole without scalar hair obtained from the Einstein-nonlinear electromagnetic theory [12, 13]. Recently, the relevant astrophysical and physical implications of the CHB were discussed in [14, 15, 16, 17].
@@
For comparison, it is desirable to involve thermodynamics of the Reissner-Nordström black hole (RNB) obtained from the Einstein-Maxwell theory but the RNB is disconnected to the CHB and the NS. Then, we have three branches for extended thermodynamic analysis. The RNB is considered as a standard form for charged black holes. Thermodynamic behavior for the CHB without the NS point is similar to that for the RNB, but its Helmholtz free energy is always positive. If the NS singular point is included as an extremal point, then the Helmholtz free energy is always negative, suggesting that the globally stable region is achieved anywhere. For the NS, its positive temperature has a maximum point and its heat capacity remains negative without having Davies point.
@@ 2 CHB and NS from the EHM theory
From $h(r)=0$, one finds two analytic horizons
$r_{+1}(m,q),\quad r_{+2}(m,q)$ (5)
whose forms are too complicated to write down here. The former is for describing the CHB while the latter is designed for the NS. These will be used to compute thermodynamic quantities and prove the first law of thermodynamics and Smarr formula. For $m=1$ case, one finds their explicit form as
$r_{+1}(1,q)$ $=$ $\frac{1}{2}\Bigg{(}1+\frac{\xi}{\sqrt{3}}+\sqrt{2-\frac{4q^{2}}{3}-\frac{1}{3}q^{4/3}(-9+8q^{2})^{1/3}+\frac{8\sqrt{3}(1-q^{2})}{4\xi}}\Bigg{)},$ (6)
$r_{+2}(1,q)$ $=$ $\frac{1}{2}\Bigg{(}1-\frac{\xi}{\sqrt{3}}+\sqrt{2-\frac{4q^{2}}{3}-\frac{1}{3}q^{4/3}(-9+8q^{2})^{1/3}-\frac{8\sqrt{3}(1-q^{2})}{4\xi}}\Bigg{)}$ (7)
@@
$\xi=\sqrt{3-2q^{2}+q^{4/3}(-9+8q^{2})^{1/3}}.$ (8)
These will be used to make pictures for the CHB and the NS.
@@
$0<\frac{q}{m}<\frac{3}{2\sqrt{2}}=1.06066(\simeq 1.06),$ (10)
which defines $q_{\rm NS}=1.06m$. For $q<q_{\rm NS}$, one finds the CHB, while for $q>q_{\rm N}$, one has the NS. The $q=q_{\rm NS}$ corresponds to the singular and extremal point.
Figure 2: Two horizons (branches) $r_{+1}(1,q)$ and $r_{+2}(1,q)$ with $m=1$ and two horizons $r_{\rm RN}(1,q)$ and $r_{\rm-RN}(1,q)$ with $m=1$ for RNB. Here, we introduce a line of $r_{+}=r_{\rm NS}(q)$ for showing the NS point and $r_{+}=r_{D}(q)$ for representing Davies line of heat capacity. For $r_{+}>r_{\rm NS}(q)$, one finds the CHB, whereas one has the NS for $r_{+}<r_{\rm NS}(q)$. The shaded region represents the positive heat capacity for the CHB, while $r_{+}>r_{\rm D}(q)$ denotes the negative heat capacity for the CHB. The NS has a negative heat capacity.
@@
which implies the condition for the existence of two horizons: $10<q/m<$ 1. Observing Fig. 2, there is no inner horizon which states that it satisfies no scalar-haired inner horizon theorem [18] and one branch is described by $r_{+1}(m,q\in[0,1.06])\geq r_{\rm RN}(m,q\in[0,1])$. Its $q$-range is beyond $[0,1]$ of RNB (overcharged), which suggests that the CHB has a scalar hair $\phi(r)$ [20, 21]. Instead, it is important to note that there exist other branch described by $r_{+2}(m,q\in[1.07,2])$. However, it is always less than $r_{\rm NS}(q)=q/\sqrt{2}$ for each $q$ and thus, it describes the NS. The blue dot denotes a point of $(q_{\rm NS}=1.06,r_{+1}(q_{\rm NS})=0.76)$ for $m=1$ which corresponds to the singular point as well as the extremal point if one takes into account the other branch $r_{+2}(m,q)$ seriously. In principle, this point will be excluded, which is distinguished from the extremal RNB of ($r_{\rm eRN}(q_{e}=1)=1$, cyan dot) for thermodynamic analysis of the RNB. In this work, we wish to consider this singular point seriously for studying the Helmholtz free energy. Finally, a yellow dot at $q=q_{D}=0.91$ represents the Davies point where heat capacity blows up.
@@ 3 Extended thermodynamic analysis
We note that the temperature $T_{\kappa}(m,q)$ together the Wald’s entropy [$S_{1}(m,q)=\pi r_{+1}^{2}(m,q)$: area law entropy after considering a non-vanishing ambiguity $\pi q^{2}/2$ with $\gamma=-2$] does not lead the first law of thermodynamics as $dM=T_{\kappa}dS_{1}-\Phi_{+1}dq$. Therefore, we have to introduce the new temperature by considering ${\cal G}=1+2\gamma X$ with $X=-(\partial\phi)^{2}/2$ as [11]
@@
To achieve the extended thermodynamics, we need to define the mass, area entropy, and new temperature for $r_{+2}(m,q)$ so that its first law and Smarr formula are satisfied. The relevant quantities for RNB are given as with the outer horizon $r_{\rm RN}(m,q)=m-\sqrt{m^{2}-q^{2}}$
@@
It is worth noting that these satisfy the first law and Smarr formula. As is shown in Fig. 3, the CHB temperature starts with $T_{1}(1,0)=1/8\pi$ (Schwarzschild temperature) and it shows a similar behavior for $q\in[0,1.06]$ as a decreasing function of the RNB. On the other hand, the positive NS temperature has a maximum point $(q=2.37,T_{2}(1,q)=1.70)$ and approaches zero as $q$ increases. This is hotter than that of the CHB.
Fig. 4 of entropy graphs shows that there exist three branches: the CHB is similar to the RNB but $S_{\rm RN}(1,q)\leq S_{1}(1,q)$, while the NS shows a quite different behavior of an increasing function of $q$. This picture is similar to the entropy picture for scalarized charged black holes obtained from the Einstein-Maxwell-scalar theory with a nonminimal coupling function $f(\Phi)=1+\alpha\Phi^{4}$, including the hot, the cold, and the RNB branches [23]. A difference is that the scalarized charged black holes do not contain any NS.
@@
Here, we allow to define the heat capacity $C_{2}(m,q)$ for the NS by replacing $r_{+1}(m,q)$ by $r_{+2}(m,q)$ in $C_{1}(m,q)$. As is shown in Fig. 2, the Davies line of $r_{D}(q)=\sqrt{5/2}q$ is derived from imposing the zero numerator of $C_{1}(1,q)$ and it splits heat capacity into negative and positive ones. As is shown in Fig. 5, the CHB heat capacity with Davies point ($q_{D}=0.91$) indicates a similar behavior as the RN heat capacity with Davies ($q_{D{\rm RN}}=0.87$) and extremal points, whereas the negative NS heat capacity has no Davies point and it decreases as $q$ increases. This implies that the CBH and RNBH can be in locally thermal equilibrium with $q$, while the NS cannot be in locally thermal equilibrium with $q$.
@@
$F_{1/2}(m,q)=\frac{r_{+1/2}(m,q)}{4}+\frac{3q^{2}}{4r_{+1/2}^{2}(m,q)}-\frac{5q^{4}}{48r_{+1/2}^{3}(m,q)}\quad F_{\rm RN}=\frac{r_{\rm RN}(m,q)}{4}+\frac{3q^{2}}{4r_{\rm RN}^{2}(m,q)}-q,$ (19)
@@
Observing Fig. 6, one finds that the CHB does not have a globally stable region when excluding the NS point as an extremal point but the NS decreases without limitation as $q$ increases. For a RNB, its free energy starts with 0.5 and arrives zero at $q=0.6$ and becomes negative. This implies that the RNB can have a globally thermodynamic stability for $0.6<q\leq 1$. If the NS point is included as an extremal point, however, the CHB has a globally stable region for whole $q\in[0,1.06]$. In this case, there may be a phase transition from the RN to the CHB because $S_{\rm RN}(1,q)\leq S_{1}(1,q)$ and $F_{1{\rm NS}}(1,q)<F_{\rm RN}(1,q)$.
@@ 4 Discussions
We have performed extended thermodynamic analysis of the CHB in $q\in[0,1.06]$ with $m=1$ and the NS in $q\in[1.07,10]$ by introducing two horizons $r_{+1/2}(m,q)$. The thermodynamic behavior of the CHB with the NS singular point is similar to that of RNB, where its temperature decreases to zero as the charge $q$ increases, its heat capacity has a Davies and extremal points, and its free energy becomes negative for $q>0.6$. In this case, a phase transition may occur from the RN to the CHB because $S_{\rm RN}(1,q)\leq S_{1}(1,q)$ and $F_{1{\rm NS}}(1,q)<F_{\rm RN}(1,q)$. However, the CHB without the NS point is quite different from the RNB, implying that the zero temperature cannot be attained and its free energy is always globally unstable. On the other hand, the NS has a maximum temperature at $q=2.37$, the heat capacity remains negative without Davies point, and its negative free energy decreases as $q$ increases.

### v2

@@ 1 Introduction
It is known that thermodynamic description of black holes is universal and can be applicable to black holes found in beyond Einstein gravity. However, it was found that the Wald’s entropy formula and /or the first law of thermodynamics does not work for the CHB [8]. To achieve the first law with the ordinary Hawking temperature, they introduced a new term $\Phi_{\phi}$ of the chemical potential and a conserved scalar charge $Q_{\phi}$. This derivation of the first law was based on the Wald’s entropy ($S_{W}=\pi(r_{+}^{2}-q^{2}/2)$) from the Noether’s theorem [9, 10] and the Hawking temperature ($T_{\kappa}=\kappa/2\pi$) obtained from the surface gravity. However, Noether’s theorem and Wald formula have ambiguities for the Einstein-Horndeski-Maxwell theory. These were fixed by deriving the new temperature and introducing the Bekenstein-Hawking entropy [11], leading to the standard first law of thermodynamics. Also, this case satisfies the Smarr formula of its integral form. We stress that this happened to the Bardeen regular black hole without scalar hair obtained from the Einstein-nonlinear electromagnetic theory [12, 13]. Recently, the relevant astrophysical and physical implications of the CHB were discussed in [14, 15, 16, 17]. The thermodynamic aspects of a charged Hordenski black hole were also studied in [18, 19].
@@
For comparison, it is desirable to involve thermodynamics of the Reissner-Nordström black hole (RNB) obtained from the Einstein-Maxwell theory but the RNB is disconnected to the CHB and the NS. Then, we have three branches for extended thermodynamic analysis. The RNB is considered as a standard form for charged black holes. Thermodynamic behavior for the CHB without the NS point is similar to that for the RNB, but its Helmholtz free energy is always positive. If the NS point is included as an extremal point, then the Helmholtz free energy is always negative, suggesting that the globally stable region is achieved anywhere. For the NS thermodyamics, its positive temperature has a maximum point and its heat capacity remains negative without having Davies point.
@@ 2 CHB and NS from the EHM theory
From $h(r)=0$, one finds four real roots as
$r_{+1}(m,q)=\frac{1}{2}\Bigg{(}m+\frac{\xi}{\sqrt{3}}+\sqrt{\chi+\frac{8\sqrt{3}m(m^{2}-q^{2})}{4\xi}}\Bigg{)},$ (5)
$r_{+2}(m,q)=\frac{1}{2}\Bigg{(}m-\frac{\xi}{\sqrt{3}}+\sqrt{\chi-\frac{8\sqrt{3}m(m^{2}-q^{2})}{4\xi}}\Bigg{)},$ (6)
$r_{+3}(m,q)=\frac{1}{2}\Bigg{(}m-\frac{\xi}{\sqrt{3}}-\sqrt{\chi-\frac{8\sqrt{3}m(m^{2}-q^{2})}{4\xi}}\Bigg{)},$ (7)
$r_{+4}(m,q)=\frac{1}{2}\Bigg{(}m+\frac{\xi}{\sqrt{3}}-\sqrt{\chi+\frac{8\sqrt{3}m(m^{2}-q^{2})}{4\xi}}\Bigg{)}$ (8)
@@
$\xi=\sqrt{3m^{2}-2q^{2}+q^{4/3}(-9m^{2}+8q^{2})^{1/3}},\quad\chi=2m^{2}-\frac{4q^{2}}{3}-\frac{1}{3}q^{4/3}(-9m^{2}+8q^{2})^{1/3}.$ (9)
For $m=1$ case, $r_{+1}(1,q)$ and $r_{+2}(1,q)$ describe the CHB and the NS, whereas $r_{+3}(1,q)$ is a negative line and $r_{4+}(1,q)$ represents a point at $q=q_{\rm NS}=3/2\sqrt{2}$. Hence, the former two (CHB and the NS) will be used to compute thermodynamic quantities and prove the first law of thermodynamics and Smarr formula.
@@
$0<\frac{q}{m}<\frac{q_{\rm NS}}{m}=\frac{3}{2\sqrt{2}m}=\frac{1.06066(\simeq 1.06)}{m}.$ (11)
For $q<q_{\rm NS}$, one finds the CHB, while for $q>q_{\rm N}$, one has the NS. The $q=q_{\rm NS}$ corresponds to the singular and extremal point because of $r_{1+}(1,q_{\rm NS})=r_{2+}(1,q_{\rm NS})=r_{4+}(1,q_{\rm NS})$.
Figure 2: Two horizons (branches) $r_{+1}(m=1,q)$ and $r_{+2}(1,q)$ and two horizons $r_{\rm RN}(1,q)$ and $r_{\rm-RN}(1,q)$ with an extremal point (cyan dot) at ($q=1,r_{\rm eRN}=1$) for RNB. Here, we introduce a line of $r_{+}=r_{\rm NS}(q)$ for showing the NS and $r_{+}=r_{D}(q)$ for representing Davies line of heat capacity. For $r_{+}>r_{\rm NS}(q)$, one finds the CHB, whereas one has the NS for $r_{+}<r_{\rm NS}(q)$. One finds the NS point (blue dot) at ($q=1.06,r_{+4}=3/4$). The shaded region represents the positive heat capacity for the CHB, while $r_{+}>r_{\rm D}(q)$ denotes the negative heat capacity for the CHB. Finally, we display $r_{+1}(m=1.1,q)$ and $r_{+2}(m=1.1,q)$.
@@
which implies a condition for the existence of two horizons: $10<q/m<$ 1.
$q$ 0 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0 1.06
$r_{+1}(1,q)$ 2 1.97981 1.954 1.917 1.867 1.802 1.719 1.61 1.46 1.23 0.84
$r_{\rm RN}(1,q)$ 2 1.9798 1.9539 1.916 1.866 1.8 1.714 1.6 1.44 1.0 N.A.
Table 1: Comparing $r_{+1}(1,q)$ with $r_{\rm RN}(1,q)$.
Observing Fig. 2, there is no inner horizon which states that it satisfies no scalar-haired inner horizon theorem [20]. We find from Table 1 that $r_{+1}(1,q\in[0,1.06])\geq r_{\rm RN}(1,q\in[0,1])$, showing no overlapping except for the $q=0$. Its $q$-range is beyond $[0,1]$ of RNB (overcharged), which suggests that the CHB has a scalar hair $\phi(r)$ [22, 23]. Instead, it is important to note that there exists the other branch described by $r_{+2}(m,q\in[1.07,2])$. However, it is always less than $r_{\rm NS}(q)=q/\sqrt{2}$ for each $q$ and thus, it describes the NS. The blue dot denotes a point of $(q_{\rm NS}=1.06,r_{+1}(q_{\rm NS})=0.75)$ for $m=1$ which corresponds to the singular point as well as the extremal point if one takes into account the other branch $r_{+2}(m,q)$ seriously. In other words, $r_{+1}(1,q)$ meets $r_{+2}(1,q)$ at the blue dot, while $r_{\rm RN}(1,q)$ meets $r_{\rm-RN}(1,q)$ at the cyan dot, leading to that two points are different. In principle, this blue point will be excluded, which is distinguished from the extremal RNB for thermodynamic analysis of the RNB. In this work, we wish to consider this blue point seriously for studying the Helmholtz free energy. Finally, a yellow dot at $q=q_{\rm D}=0.91$ represents the Davies point where heat capacity blows up. For $m\not=1$, we display $r_{+1}(m=1.1,q)$ and $r_{+2}(m=1.1,q)$, showing qualitatively similar behaviors as $r_{+1}(1,q)$ and $r_{+2}(1,q)$.
@@ 3 Extended thermodynamic analysis
We note that the temperature $T_{\kappa}(m,q)$ together the Wald’s entropy [$S_{1}(m,q)=\pi r_{+1}^{2}(m,q)$: area law entropy after considering a non-vanishing ambiguity $\pi q^{2}/2$ with $\gamma=-2$] does not lead the first law of thermodynamics as $dM=T_{\kappa}dS_{1}+\Phi_{+1}dq$. Therefore, we have to introduce the new temperature by considering ${\cal G}=1+2\gamma X$ with $X=-(\partial\phi)^{2}/2$ as [11]
@@
To achieve the extended thermodynamics, we need to define the mass, area entropy, and new temperature for $r_{+2}(m,q)$ so that its first law and Smarr formula are satisfied. The relevant quantities for RNB are given as with the outer horizon $r_{\rm RN}(m,q)=m+\sqrt{m^{2}-q^{2}}$
@@
It is worth noting that these satisfy the first law and Smarr formula. As is shown in Fig. 3, the CHB temperature starts with $T_{1}(1,0)=1/8\pi$ (Schwarzschild temperature) and it shows a similar behavior for $q\in[0,1.06]$ as a decreasing function of the RNB. On the other hand, the positive NS temperature has a maximum point $(q=2.37,T_{2}(1,q)=1.70)$ and approaches zero as $q$ increases. This is hotter than that of the CHB. We note that the CHB and RNB extremal temperatures (blue and red dots) show zero temperatures.
Fig. 4 of entropy graphs shows that there exist three branches: the CHB is similar to the RNB but $S_{\rm RN}(1,q)\leq S_{1}(1,q)$, while the NS shows a quite different behavior of an increasing function of $q$. The blue and cyan points represent different entropies for the extremal CHB and RNB, respectively. This picture is similar to the entropy picture for scalarized charged black holes obtained from the Einstein-Maxwell-scalar theory with a nonminimal coupling function $f(\Phi)=1+\alpha\Phi^{4}$, including the hot, the cold, and the RNB branches [25]. A difference is that the scalarized charged black holes do not contain the NS point and NS.
@@
Here, we allow to define the heat capacity $C_{2}(m,q)$ for the NS by replacing $r_{+1}(m,q)$ by $r_{+2}(m,q)$ in $C_{1}(m,q)$. As is shown in Fig. 2, the Davies line of $r_{D}(q)=\sqrt{5/2}q$ is derived from imposing the zero numerator of $C_{1}(1,q)$ and it splits heat capacity into negative and positive ones. As is shown in Fig. 5, the CHB heat capacity with Davies point ($q_{D}=0.91$) indicates a similar behavior as the RN heat capacity with Davies ($q_{D{\rm RN}}=0.87$) and extremal points, whereas the negative NS heat capacity has no Davies point and it decreases as $q$ increases. This implies that the CBH and RNB can be in locally thermal equilibrium with $q$, while the NS cannot be in locally thermal equilibrium with $q$.
@@
$F_{1/2}(m,q)=\frac{r_{+1/2}(m,q)}{4}+\frac{3q^{2}}{4r_{+1/2}(m,q)}-\frac{5q^{4}}{48r_{+1/2}^{3}(m,q)}\quad F_{\rm RN}=\frac{r_{\rm RN}(m,q)}{4}+\frac{3q^{2}}{4r_{\rm RN}(m,q)}-q,$ (20)
@@
Observing Fig. 6, one finds that the CHB does not have a globally stable region when excluding the NS point as an extremal point but the NS decreases without limitation as $q$ increases. For a RNB, its free energy starts with 0.5 and arrives zero at $q=0.6$ and becomes negative. This implies that the RNB can have a globally thermodynamic stability for $0.6<q\leq 1$. If the NS point is included as an extremal point, however, the CHB has a globally stable region for whole $q\in[0,1.06]$. In this case, there may be a phase transition from the RN to the CHB because $S_{\rm RN}(1,q)\leq S_{\rm 1NS}(1,q)$ and $F_{\rm 1NS}(1,q)<F_{\rm RN}(1,q)$.
@@ 4 Discussions
We have performed extended thermodynamic analysis of the CHB in $q\in[0,1.06]$ with $m=1$ and the NS in $q\in[1.07,10]$ by introducing two horizons $r_{+1/2}(m,q)$. We note that this CHB and NS with scalar hair has no inner horizon while the RNB has outer and inner horizons. Let us compare the thermodynamic functions for CHB with those for RNB.
The thermodynamic behavior of the CHB with the NS point is similar to that of RNB in $q\in[0,1]$, where their temperatures decrease to zero as the charge $q$ increases, their entropies decreases as $q$ increases with terminations, their heat capacities have a Davies and extremal points, and their free energies have negative regions. In this case, a phase transition from the RNB to the CHB may occur because $S_{\rm RN}(1,q)\leq S_{\rm 1NS}(1,q)$ and $F_{\rm 1NS}(1,q)<F_{\rm RN}(1,q)$. However, there is no overlapping between CHB and RNB (see Table1) because their horizon structures are different (see Fig. 1) and thus, the only correspondence between CHB and RNB arises from the $q=0$ limit, leading to the Schwarzschild black hole. This implies that thermodynamic values of $q=0$ CHB are those of $q=0$ RNB, leading to thermodynamic values of Schwarzschild BH.
On the other hand, the CHB without the NS point is quite different from the RNB, implying that the zero temperature cannot be attained and its free energy is always globally unstable. Furthermore, the NS thermodynamics may imply that its temperature has a maximum point at $q=2.37$, its heat capacity remains negative without Davies point, and its negative free energy decreases as $q$ increases without limitation on $q$. However, we are concerned that the NS inside the NS point has a horizon to define its thermodynamics.